OpenStudy (anonymous):

lim θ→0 sin(cos 8θ)/sec 5θ

OpenStudy (anonymous):

lim θ→0 sin(cos 8θ)/cos8θ . cos 8θ/sec5θ then I get cos 8θ. cos 5θ but i don't know how to continue it

myininaya (myininaya):

\[\lim_{\theta \rightarrow 0}\frac{ \sin (\cos(8 \cdot \theta))}{\sec(5 \cdot \theta)}\] Why don't you just use direct substitution :)

myininaya (myininaya):

\[\sec( 5 \cdot 0)=\sec(0)=1\]

myininaya (myininaya):

Bottom is not zero so we are good

OpenStudy (anonymous):

cos 8θ=0 right? is sin 1 is 1?

OpenStudy (anonymous):

i tried 1 for the answer but it is wrong

myininaya (myininaya):

\[\sin( \cos(8 \cdot 0))=\sin(\cos(0))=\sin(1)\] We have sin(1)/1=sin(1) We are done No sin(1) totally does not =1

OpenStudy (anonymous):

thankkksss you are right, I gave you medal already

myininaya (myininaya):

lol ok thanks for the medal you are sweet